Find 𝑷(𝟐) and the value of π’Œ in the following discrete probability distribution. Show your complete solution


x=0,1,2,3,4

p(x)=0.13,0.27,π’Œ,0.09,0.12


(a) find π’Œ

(b) 𝑷(𝟐)


1
Expert's answer
2022-03-02T11:08:21-0500

(a)


"\sum _ip(x_i)=0.13+0.27+k+0.09+0.12=1"

"k=0.39"

(b)

"P(2)=0.39"